If an object is thrown with a velocity of v feet per second at an angle of θ with the horizontal, then its flight can be modeled by,

x = (v cos θ ) t and y = v (sin θ ) t - 16 t2 + h

where t is in seconds and h is the object's initial height in feet above the ground.

x is the horizontal position and y is the vertical position, and - 16 t2 represents gravity pulling on the object.

Depending on the units involved in the problem, use g = 32 ft/ s2 or g 9.8 m/ s2.

Assume that the ball is hit with an initial velocity of 140 feet per second at an angle of  45° to the horizontal, making contact 3 feet above the ground.

  1. Find the parametric equations to model the path of the baseball.
  2. Where is the ball after 2 seconds?
  3. How long is the ball in the air?
  4. Is it a home run?
Answer #1

л - қ9) + у 1 = u(sin O) +- 16 t²th v= 140 pearls 2 9 = 45° и: т у substituting this in Мла мал май ол (140 и 4) t - it * * иа) t = 6. а11 - О. oъб 16 2 since оа &ация in time period cannot be negarine, you the air 6.2 1949 Орлоо ) 40 Sa x 6.2 13 615